OpenStudy (anonymous):

The sides of a square are five to the power of two fifths inches long. What is the area of the square?

OpenStudy (anonymous):

dont you multiply @Bill_Cipher

OpenStudy (anonymous):

i think the answer is 25^4/45

OpenStudy (anonymous):

if you mean area of a square then yes you multiply the base of the square by the length, but since the base and length on a square are the same you square

OpenStudy (anonymous):

Remember [5^(2/5)]^2=5^(2/5*2)

OpenStudy (anonymous):

wat? it says find the area of the square

OpenStudy (anonymous):

i am so confused

OpenStudy (anonymous):

what are you confused on?

OpenStudy (anonymous):

wait is my answer wrong

OpenStudy (anonymous):

25^(4/45) isn't the area of the square

OpenStudy (anonymous):

the answer is 5^(2/5*2)

OpenStudy (anonymous):

what does this simplify to?

OpenStudy (anonymous):

5^4 over 25 ?

OpenStudy (anonymous):

2/5*2/1=4/5 so we have 5^(4/5)

OpenStudy (anonymous):

ohhhhhhh ok I think i get it that you <3

OpenStudy (anonymous):

do you need the answer in radical form?

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

could you help me with another

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which of the following statements is false? The sum of two rational numbers is always rational. The product of a nonzero rational number and an irrational number is always irrational. The product of two rational numbers is always rational. The sum of two irrational numbers is always rational.

OpenStudy (anonymous):

the last one. Example that disproves it: pi+pi=2pi 2pi isn't rational

OpenStudy (anonymous):

oh ok thank you I own you big time hate math lol

OpenStudy (anonymous):

your welcome. Was that your last question?

OpenStudy (anonymous):

yeah i think so thank you so much